Let G be the group of integers under the operation of addition,

and let H = {3k | k ∈ Z}. Is H a subgroup of G


1
Expert's answer
2022-04-18T00:10:56-0400

Since "G" is a group and "H" is a subset of "G" (all elements of "H" are integers), it is enough to check that for any two elements "a,b\in H" we have: "a+b\in H". It is the subgroup criterion. Suppose that "a,b\in H". It means that they can be presented as: "a=3k,b=3z," "k,z\in{\mathbb{Z}}". Then, "a+b=3k+3z=3(k+z)". As we can see from the form, "a+b\in H." Thus, "H" is a subgroup of "G."